OpenStudy (anonymous):

Explain how you get the answer PLease, i know how to get it and i know the process i just keep getting the wrong answer

OpenStudy (anonymous):

The Midpoint UV is (5, -11). The coordinates of one endpoint are U(3,5). What is the coordinate of the endpoin V.

Nnesha (nnesha):

the formula for midpoint is \[\huge\rm (\frac{ x_1+x_2 }{ 2 } , \frac{ y_1+y_2 }{ 2 })\]given is (x_1 ,y_1) order pair and midpoint (x,y) we can separate both x's and y's \[\large\rm \frac{ x_1+x_2 }{ 2 }=(x-coordinate ~of~midpoint)\] \[\large\rm \frac{ y_1 +y_2 }{ 2 }=(y-coordinate ~of ~midpoint)\] replace (x_1 ,y_1) with the given end point

OpenStudy (anonymous):

So would x_1 and y_1 be the midpoint and x_2 and y_2 the endpoint or the other way around?

Nnesha (nnesha):

no (x_1 ,y_1) is the endpoint (3,5)

Nnesha (nnesha):

you have to solve for x_2 and y_2

OpenStudy (anonymous):

Oh! Okay

Nnesha (nnesha):

like to find x-coordinate of 2nd endpoint \[\large\rm \frac{ x_1+x_2 }{ 2 }=(x-coordinate ~of~midpoint)\] \[\large\rm \frac{ 3+x_2 }{ 2 }=5\] now solve for x_2

OpenStudy (anonymous):

7?

Nnesha (nnesha):

yes right you can check your work \[\frac{ 3+7 }{ 2 }\] if you get 5 then yes ofc it's right bec the midpoint (5,-11) is given

OpenStudy (anonymous):

so what would the other endpoint be?

Nnesha (nnesha):

we got the x-coordinate now find y-coordinate

Nnesha (nnesha):

\[\large\rm \frac{ y_1 +y_2 }{ 2 }=(y-coordinate ~of ~midpoint)\]

OpenStudy (anonymous):

5+y_2/2=-11?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

-27?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

so 7, -27?

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Thank you so much!

Nnesha (nnesha):

np :=) just an advice the drawing in ur pfp is really good you should post that in design if youdrw that pic

OpenStudy (anonymous):

Thank you hahah